# QuantFactor REINFORCE: Mining Steady Formulaic Alpha Factors with Variance-bounded REINFORCE

### v2

@@ V-F Ablation Study
Figure 9: Ablation experiments in the constituent stocks of CSI300, CSI500, CSI1000, SPX, DJI, and NDX, where the baseline and reward shaping are removed separately. All the curves are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
To investigate the role of the two improvements of QFR, we designed two variants including: QFR without baseline represents that the policy network is updated directly using REINFORCE, and utilizes the reward function defined in (11); QFR without reward shaping represents that IC is used directly as reward without any IR test. Fig. 9 presents the ablation results of the two improvements. The QFR can provide prominent performance among all six index constituent stocks, indicating that all improvements play important roles. When removing the baseline leads to a degradation in the quality of the factors, it implies that the baseline can help effectively reduce the variance in the training process and achieve better performance. While QFR without reward shaping learns faster during the beginning learning stage, it finally achieves worse performance, which may be due to an overemphasis on the absolute returns of factors while neglecting their volatility resistance, leading to a local optimum trap. In summary, our proposed QFR conditioned on all improvements gives the best performance, indicating that baseline and reward shaping are complementary parts of QFR.
@@ VI Conclusion
In this paper, we have proposed a novel reinforcement learning algorithm, QuantFactor REINFORCE (QFR), for mining formulaic alpha factors. QFR leverages the advantages of discarding the critic network while theoretically addressing its limitations of high variance by introducing a greedy baseline. Additionally, the incorporation of IR as a reward shaping mechanism encourages the generation of stable alpha factors that can better adapt to changing market conditions.
Our extensive experiments on real-world asset datasets demonstrate the superiority of QFR over existing factor mining methods. QFR generates alpha factors with higher correlation to asset returns and stronger ability to generate excess profits. It also achieves better performance compared to other state-of-the-art reinforcement learning algorithm when mining formulaic alpha factors.The experimental results also align closely with the theoretical analysis, validating the effectiveness of our proposed algorithm.
We conclude that QFR is a promising approach for mining formulaic alpha factors. Its interpretability, stability, and efficiency make it a valuable tool for quantitative finance applications. Future research directions include exploring more sophisticated reward shaping mechanisms and applying QFR to other financial tasks such as portfolio optimization and risk management.

### v3

@@ V-F Ablation Study
To investigate the role of the two improvements of QFR, we designed two variants including: QFR without baseline represents that the policy network is updated directly using REINFORCE; QFR without reward shaping represents that IC is used directly as reward without any IR test. Fig. 9 presents the ablation results. QFR can provide prominent performance among all six index constituent stocks, indicating that all improvements play important roles. When removing the baseline leads to a degradation in the quality of the factors, it implies that the baseline can help reduce the variance. While QFR without reward shaping learns faster at the beginning, it finally falls to local optima, which may be due to an overemphasis on the absolute returns. In summary, QFR conditioned on all improvements gives the best performance, indicating that baseline and reward shaping are complementary parts of QFR.
@@ VI Conclusion
In this paper, we have proposed a novel RL algorithm, QuantFactor REINFORCE (QFR), for mining formulaic alpha factors. QFR leverages the advantages of discarding the critic network while theoretically addressing its limitations of high variance by introducing a greedy baseline. Additionally, the incorporation of IR as a reward shaping mechanism encourages the generation of stable alpha factors that can better adapt to changing market conditions. Our extensive experiments on real-world asset datasets demonstrate QFR achieves better performance compared to other state-of-the-art RL algorithm when mining formulaic alpha factors. It also generates superior alpha factors over existing factor mining methods. We conclude that QFR is a promising approach for mining formulaic alpha factors. Beyond financial markets, the symbolic regression capabilities of QFR could be extended to diverse signal processing tasks that require interpretable, formula-based models, such as symbolic regression for time-frequency analysis[64], feature extraction in image denoising[65], and adaptive filtering in speech processing[66]. Future research directions include integrating matrix tri-factorization techniques to enhance QFR’s capability in capturing cross-stock correlations and leveraging community detection techniques to uncover latent asset clusters and investor behavior patterns.